OpenStudy (anonymous):

how do i integrate (1+r)^-2 dr?

OpenStudy (anonymous):

\[\int \frac{1}{(1+r)^2}dt\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the end is dr not dt

OpenStudy (anonymous):

use the power rule backwards

OpenStudy (anonymous):

how do i use the power rule backwards?

OpenStudy (anonymous):

\[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

can x in that case be a function, like 1+r?

OpenStudy (anonymous):

so \[\int (1+r)^{-2}dr=\frac{(1+r)^{-1}}{-1}\]

OpenStudy (anonymous):

better written as \(-\frac{1}{r+1}\)

OpenStudy (anonymous):

yes , you can treat \(1+r\) as \(x\) because it is just \(1+r\) and not say \(1+2r\)

OpenStudy (anonymous):

um how are 1+r and 1+2r different when it comes to using the powerrule for integrals?

OpenStudy (anonymous):

\[\int\limits f^n \left( x \right)f'\left( x \right)dx=\frac{ f ^{n+1}\left( x \right) }{ x+1 }\]

OpenStudy (anonymous):

i thought (1+r) was a function itself.

OpenStudy (anonymous):

is there a backwards chain rule?

OpenStudy (anonymous):

Let \(x=1+r\) and then \(dx = dr\)

OpenStudy (anonymous):

okay, what is next?

OpenStudy (anonymous):

\[ \int x^{-2}\;dx \]This is a simple integral.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i have -x^-1 do i replace x with 1+r before evaluating the defined intgeral?

OpenStudy (anonymous):

sure